# A generalization of the relaxation-based matrix splitting iterative method for solving the system of generalized absolute value equations

### v1

@@ 3 Convergence analysis
$a(|1-\theta|+\theta d)>1\quad\text{and}\quad a+|1-\theta|+\theta d>2.$ (3.36)
In addition, it follows from (3.3) that
$\theta\alpha c<-2c\beta+1-(a+|1-\theta|+\theta d)+a(|1-\theta|+\theta d).$ (3.37)
On the other hand, from the condition (3.2), we obtain
$-2(a|1-\theta|+\theta ad+1)<-2c\beta<2(1-a|1-\theta|-\theta ad).$ (3.38)
Hence, it follows from (3.36), (3.37) and (3.38) that
@@
which is a contradiction. Hence, if (3.2) and (3.3) hold, we have $a<1$ and $|1-\theta|+\theta d<1$.
Then it follows from (3.35) that
$\|y^{*}-\bar{y}^{*}\|\leq\dfrac{\theta\alpha+2\beta}{1-|1-\theta|-\theta d}\|x^{*}-\bar{x}^{*}\|.$ (3.39)
According to (3.39), (3.34) and (3.3), we obtain
@@
Remark 3.1.
In Theorem 3.1, we assume that $Q$ is nonsingular so that $y^{*}=Q^{-1}|x^{*}|$ is obtained. If we are not interested in the limit of the auxiliary sequence $\{y^{(k)}\}_{k=0}^{\infty}$, the nonsingularity hypothesis of $Q$ is not required. In this case, the results of Theorem 3.1 except that of $\{y^{(k)}\}_{k=0}^{\infty}$ still hold. For the sake of clarity, we always remain the nonsingularity of $Q$ in the following.
Based on Theorem 3.1, we have the following Corollary 3.1.
Corollary 3.1.
Assume that $A,B,H\in\mathbb{R}^{n\times n}$ and $Q\in\mathbb{R}^{n\times n}$ is nonsingular. Let $A=M-N$ and $Q=Q_{1}-Q_{2}$ with $M$ and $Q_{1}$ being nonsingular. Let $a,c,d,\alpha$ and $\beta$ be defined as in (3.1). If
$a<1,~{}c\beta<ad+1,~{}c(2\beta+\alpha)<(1-a)(1-d),$ (3.40)
@@
$\dfrac{2c\beta}{(1-a)(1-d)-c\alpha}<\theta<\dfrac{2(1-a)-2c\beta}{(1-a)(1+d)+c\alpha},$ (3.41)
then the sequence $\{(x^{(k)},~{}y^{(k)})\}^{\infty}_{k=0}$ generated by (2.3) converges to $(x^{*},~{}Q^{-1}|x^{*}|)$ and $x^{*}$ is the unique solution of the GAVEs (1.1).
@@
The corollary will be proved if we can show that (3.40) and (3.41) imply (3.2) and (3.3). The proof is divided into two cases.
@@
    We first consider $\dfrac{2c\beta}{(1-a)(1-d)-c\alpha}<\theta\leq 1$. It follows from (3.40) that
@@
    $\dfrac{c\beta-(a-1)}{a(d-1)}<0<\theta,$ (3.42)
    $\theta\leq 1<\dfrac{c\beta-(a+1)}{a(d-1)}.$ (3.43)
    From the inequality (3.42), we have
@@
    $a(1-\theta)+\theta ad-c\beta<1.$ (3.44)
    From the inequality (3.43), we have
@@
    $a(1-\theta)+\theta ad-c\beta>-1.$ (3.45)
    Note that $0<\theta\leq 1$, combining (3.44) and (3.45), we have
@@
    which is the inequality (3.2).
    On the other hand, it follows from (3.40) that $c\alpha<(1-a)(1-d)$ (otherwise, we get $02c\beta<$ 0, a contradiction). Multiplying both sides of $\dfrac{2c\beta}{(1-a)(1-d)-c\alpha}<\theta$ by $(1-a)(1-d)-c\alpha$, we have
    $2c\beta<\theta(1-a)(1-d)-\theta c\alpha.$ (3.46)
    From (3.46), we infer that (remind that $0<\theta\leq 1$)
@@
    Now we turn to $1<\theta<\dfrac{2(1-a)-2c\beta}{(1-a)(1+d)+c\alpha}$. On the condition of (3.40) and (3.41), we have
    $\dfrac{c\beta+a-1}{a(d+1)}<1$ $<\theta,$ (3.47)
    $\theta<\dfrac{2(1-a)-2c\beta}{(1-a)(1+d)+c\alpha}$ $<\dfrac{c\beta+a+1}{a(d+1)}.$ (3.48)
    From (3.47), we have
@@
    $a(\theta-1)+\theta ad-c\beta>-1.$ (3.49)
    From (3.48), we have
@@
    $a(\theta-1)+\theta ad-c\beta<1.$ (3.50)
    Note that $\theta>1$, combining (3.49) and (3.50), we have
@@
    which is the inequality (3.2).
    On the other hand, if $a<1$, we have $(1-a)(1+d)+c\alpha>0$. According to the first inequality of (3.48), we have
@@
In the following, we will apply the results of the Theorem 3.1 to some existing methods.
@@
$\tau_{\rm RMS}\|M_{\rm RMS}^{-1}B\|<(1-\|M_{\rm RMS}^{-1}N_{\rm RMS}\|)(1-|1-\tau_{\rm RMS}|),$ (3.52)
then the GAVEs (1.1) has a unique solution $x^{*}$ for any $b\in\mathbb{R}^{n}$ and the sequence $\{x^{(k)}\}_{k=0}^{\infty}$ generated by the RMS iteration (1.3) converges to the $x^{*}$.
Remark 3.2.
The conditions (3.51) and (3.52) appear in [26, Lemma 3]. However, the existence and uniqueness of the solution to the GAVEs (1.1) is not considered there.
@@
$\|{(M_{\rm NMS}+\Omega_{\rm NMS})}^{-1}(N_{\rm NMS}+\Omega_{\rm NMS})\|+\|{(M_{\rm NMS}+\Omega_{\rm NMS})}^{-1}B\|<1,$ (3.53)
then the GAVEs (1.1) has a unique solution $x^{*}$ for any $b\in\mathbb{R}^{n}$ and the sequence $\{x^{(k)}\}_{k=0}^{\infty}$ generated by the NMS iteration [33] converges to the $x^{*}$.
Remark 3.3.
In [33, Corollary 4.2.], the authors show that the sequence $\{x^{(k)}\}_{k=0}^{\infty}$ generated by the NMS iteration converges to a solution of the GAVEs (1.1) if
$\|(M_{\rm NMS}+\Omega_{\rm NMS})^{-1}\|(\|N_{\rm NMS}+\Omega_{\rm NMS}\|+\|B\|)<1,$ (3.54)
which implies (3.53) while the reverse is generally not true. Moreover, the existence and uniqueness of the solution to the GAVEs (1.1) is not considered in [33, Corollary 4.2.].
Corollary 3.4.
@@
$\|(A+\Omega_{\rm MN})^{-1}\Omega_{\rm MN}\|+\|(A+\Omega_{\rm MN})^{-1}B\|<1,$ (3.55)
then the GAVEs (1.1) has a unique solution $x^{*}$ for any $b\in\mathbb{R}^{n}$ and the sequence $\{x^{(k)}\}_{k=0}^{\infty}$ generated by the MN iteration [27] converges to the $x^{*}$.
Remark 3.4.
In [27, Theorem 3.1], the authors show that the sequence $\{x^{(k)}\}_{k=0}^{\infty}$ generated by the MN iteration converges to a solution of the GAVEs (1.1) if
$\|(A+\Omega_{\rm MN})^{-1}\|(\|\Omega_{\rm MN}\|+\|B\|)<1.$ (3.56)
It is obvious that (3.56) implies (3.55) while the reverse is generally not true. In addition, the existence and uniqueness of the solution to the GAVEs (1.1) is not considered in [27, Theorem 3.1].
Corollary 3.5.
Assume that $A,B\in\mathbb{R}^{n\times n}$ are given and $A$ is nonsingular. If $\|A^{-1}B\|<1$, then the GAVEs (1.1) has a unique solution $x^{*}$ for any $b\in\mathbb{R}^{n}$ and the sequence $\{x^{(k)}\}_{k=0}^{\infty}$ generated by the Picard iteration [25] converges to the $x^{*}$.
Remark 3.5.
@@
$|1-\alpha_{\rm MGSOR}||1-\beta_{\rm MGSOR}|<1$
@@
$\alpha_{\rm MGSOR}\beta_{\rm MGSOR}\|A^{-1}BQ_{\rm MGSOR}\|\|Q_{\rm MGSOR}^{-1}\|<(1-|1-\alpha_{\rm MGSOR}|)(1-|1-\beta_{\rm MGSOR}|),$
then the GAVEs (1.1) has a unique solution $x^{*}$ for any $b\in\mathbb{R}^{n}$ and the sequence $\{(x^{(k)},y^{(k)})\}_{k=0}^{\infty}$ generated by the MGSOR iteration (2.9) converges to $(x^{*},y^{*}=Q_{\rm MGSOR}^{-1}|x^{*}|)$ for any given initial vectors $x^{(0)},y^{(0)}\in\mathbb{R}^{n}$.
When $B=I$ we can have the following Corollary 3.7 directly from Corollary 3.6.
Corollary 3.7.
@@
$|1-\alpha_{\rm MGSOR}||1-\beta_{\rm MGSOR}|<1$ (3.57)
@@
$\alpha_{\rm MGSOR}\beta_{\rm MGSOR}\|A^{-1}Q_{\rm MGSOR}\|\|Q_{\rm MGSOR}^{-1}\|<(1-|1-\alpha_{\rm MGSOR}|)(1-|1-\beta_{\rm MGSOR}|),$ (3.58)
then the AVEs (1.2) has a unique solution $x^{*}$ for any $b\in\mathbb{R}^{n}$ and the sequence $\{(x^{(k)},y^{(k)})\}_{k=0}^{\infty}$ generated by the MGSOR iteration (2.5) converges to $(x^{*},y^{*}=Q_{\rm MGSOR}^{-1}|x^{*}|)$ for any given initial vectors $x^{(0)},y^{(0)}\in\mathbb{R}^{n}$.
Remark 3.6.
The convergence conditions (3.57) and (3.58) are new, which are not appeared in [31]. In addition, the existence and uniqueness of the solution to the AVEs (1.2) is not considered in [31].
When $\alpha_{\rm MGSOR}=1$, $\beta_{\rm MGSOR}=\tau_{\rm MFPI}$ and $Q_{\rm MGSOR}=Q_{\rm MFPI}$, the following Corollary 3.8 follows from Corollary 3.6.
Corollary 3.8.
@@
then the GAVEs (1.1) has a unique solution $x^{*}$ for any $b\in\mathbb{R}^{n}$ and the sequence $\{(x^{(k)},y^{(k)})\}_{k=0}^{\infty}$ generated by the MFPI iteration (2.11) converges to $(x^{*},y^{*}=Q_{\rm MFPI}^{-1}|x^{*}|)$ for any given initial vectors $x^{(0)},y^{(0)}\in\mathbb{R}^{n}$.
When $\alpha_{\rm MGSOR}=1$, $\beta_{\rm MGSOR}=\tau_{\rm MFPI}$, $Q_{\rm MGSOR}=Q_{\rm MFPI}$ and $B=I$ we can establish the following Corollary 3.9 from Corollary 3.7.
Corollary 3.9.
@@
[…]
then the AVEs (1.2) has a unique solution $x^{*}$ for any $b\in\mathbb{R}^{n}$ and the sequence $\{(x^{(k)},y^{(k)})\}_{k=0}^{\infty}$ generated by the MFPI iteration (1.10) converges to $(x^{*},y^{*}=Q_{\rm MFPI}^{-1}|x^{*}|)$ for any given initial vectors $x^{(0)},y^{(0)}\in\mathbb{R}^{n}$.
Remark 3.7.
In [29], the authors show that the sequence $\{x^{(k)}\}_{k=0}^{\infty}$ generated by the MFPI iteration (1.10) converges to the unique solution $x^{*}$ of the AVEs (1.2) if
[…]
@@
In addition, $0<\nu<\dfrac{1}{\ell\sqrt{1+\varrho^{2}}}$ implies $\dfrac{1-\sqrt{1-\ell^{2}\nu^{2}}}{1-\ell\varrho\nu}>0$ and $\frac{2}{1+\|A^{-1}Q_{\rm MFPI}\|\|Q^{-1}_{\rm MFPI}\|}\geq\frac{2}{1+\|A^{-1}\|\|Q_{\rm MFPI}\|\|Q^{-1}_{\rm MFPI}\|}>\dfrac{1+\sqrt{1-\ell^{2}\nu^{2}}}{1+\ell\varrho\nu}$. Hence, (3.59) is weaker than (3.60).
If $Q_{\rm MFPI}=I$ and $\tau_{\rm MFPI}=\tau_{\rm FPI}$, the following corollary follows from Corollary 3.8.
Corollary 3.10.
@@
then the GAVEs (1.1) has a unique solution $x^{*}$ for any $b\in\mathbb{R}^{n}$ and the sequence $\{(x^{(k)},y^{(k)})\}_{k=0}^{\infty}$ generated by the FPI iteration (2.8) converges to $(x^{*},y^{*}=|x^{*}|)$ for any given initial vectors $x^{(0)},y^{(0)}\in\mathbb{R}^{n}$.
When $Q_{\rm MFPI}=I$, $\tau_{\rm MFPI}=\tau_{\rm FPI}$ and $B=I$ we can have the following result follows from Corollary 3.9.
Corollary 3.11.
@@
[…]
then the AVEs (1.2) has a unique solution $x^{*}$ for any $b\in\mathbb{R}^{n}$ and the sequence $\{(x^{(k)},y^{(k)})\}_{k=0}^{\infty}$ generated by the FPI iteration (1.8) converges to $(x^{*},y^{*}=|x^{*}|)$ for any given initial vectors $x^{(0)},y^{(0)}\in\mathbb{R}^{n}$.
Remark 3.8.
In [10], Ke show that the sequence $\{x^{(k)}\}_{k=0}^{\infty}$ generated by the FPI iteration (1.8) converges to the unique solution of the AVEs (1.2) if
[…]
It is obvious that $0<\frac{1-\sqrt{1-\nu^{2}}}{1-\nu}$ and $\dfrac{1+\sqrt{1-\nu^{2}}}{1+\nu}<\dfrac{2}{1+\nu}$ when $10<\nu<$ 1. Hence, $\{\nu:0<\nu<\frac{\sqrt{2}}{2}\}\subseteq\{\nu:0<\nu<1\}$ and $\{\tau_{\rm FPI}:\frac{1-\sqrt{1-\nu^{2}}}{1-\nu}<\tau_{\rm FPI}<\frac{1+\sqrt{1-\nu^{2}}}{1+\nu}\}\subseteq\{\tau_{\rm FPI}:0<\tau_{\rm FPI}<\frac{2}{1+\nu}\}$. In other words, (3.61) is weaker than (3.62). Moreover, in the sense of (3.61), the modification strategy of the FPI iteration [29] seems redundant in theory.
When $\alpha_{\rm MGSOR}=\beta_{\rm MGSOR}=\omega_{\rm SOR}$ and $Q_{\rm MGSOR}=I$, the following corollary follows from Corollary 3.6.
Corollary 3.12.
@@
[…]
then the GAVEs (1.1) has a unique solution $x^{*}$ for any $b\in\mathbb{R}^{n}$ and the sequence $\{(x^{(k)},y^{(k)})\}_{k=0}^{\infty}$ generated by the SOR-like iteration (2.10) converges to $(x^{*},y^{*}=|x^{*}|)$ for any given initial vectors $x^{(0)},y^{(0)}\in\mathbb{R}^{n}$.
When $\alpha_{\rm MGSOR}=\beta_{\rm MGSOR}=\omega_{\rm SOR}$, $Q_{\rm MGSOR}=I$ and $B=I$ we can have the following result from Corollary 3.7.
Corollary 3.13.
@@
[…]
then the AVEs (1.2) has a unique solution $x^{*}$ for any $b\in\mathbb{R}^{n}$ and the sequence $\{(x^{(k)},y^{(k)})\}_{k=0}^{\infty}$ generated by the SOR-like iteration (1.7) converges to $(x^{*},y^{*}=|x^{*}|)$ for any given initial vectors $x^{(0)},y^{(0)}\in\mathbb{R}^{n}$.
Remark 3.9.
In [9, Corollary 3.1], under a kernel norm, Ke and Ma show that the SOR-like iteration converges to the unique solution of the GAVEs (1.1) if
[…]
If $0<\nu<\left(\frac{1+\sqrt{5}}{5+\sqrt{5}}\right)^{2}$, we have $\frac{2-2\sqrt{\nu}}{1-\nu}>1+\frac{2}{3+\sqrt{5}}\geq\min\left\{1+\frac{2}{3+\sqrt{5}},\sqrt{\frac{2}{(3+\sqrt{5})\nu}}\right\}$, which implies that (3.63) is weaker than (3.64).

### v2

@@ 3 Convergence analysis
$a(|1-\theta|+\theta d)>1\quad\text{and}\quad a+|1-\theta|+\theta d>2.$ (42)
In addition, it follows from (27) that
$\theta\alpha c<-2c\beta+1-(a+|1-\theta|+\theta d)+a(|1-\theta|+\theta d).$ (43)
On the other hand, from the condition (26), we obtain
$-2(a|1-\theta|+\theta ad+1)<-2c\beta<2(1-a|1-\theta|-\theta ad).$ (44)
Hence, it follows from (42), (43) and (44) that
@@
which is a contradiction. Hence, if (26) and (27) hold, we have $a<1$ and $|1-\theta|+\theta d<1$.
Then it follows from (41) that
$\|y^{*}-\bar{y}^{*}\|\leq\dfrac{\theta\alpha+2\beta}{1-|1-\theta|-\theta d}\|x^{*}-\bar{x}^{*}\|.$ (45)
According to (45), (40) and (27), we obtain
@@
Remark 1.
In Theorem 1, we assume that $Q$ is nonsingular so that $y^{*}=Q^{-1}|x^{*}|$ is obtained. If we are not interested in the limit of the auxiliary sequence $\{y^{(k)}\}_{k=0}^{\infty}$, the nonsingularity hypothesis of $Q$ is not required. In this case, the results of Theorem 1 except that of $\{y^{(k)}\}_{k=0}^{\infty}$ still hold. For the sake of clarity, we always remain the nonsingularity of $Q$ in the following.
Based on Theorem 1, we have the following Corollary 1.
Corollary 1.
Assume that $A,B,H\in\mathbb{R}^{n\times n}$ and $Q\in\mathbb{R}^{n\times n}$ is nonsingular. Let $A=M-N$ and $Q=Q_{1}-Q_{2}$ with $M$ and $Q_{1}$ being nonsingular. Let $a,c,d,\alpha$ and $\beta$ be defined as in (25). If
$a<1,~{}c\beta<ad+1,~{}c(2\beta+\alpha)<(1-a)(1-d),$ (46)
@@
$\dfrac{2c\beta}{(1-a)(1-d)-c\alpha}<\theta<\dfrac{2(1-a)-2c\beta}{(1-a)(1+d)+c\alpha},$ (47)
then the sequence $\{(x^{(k)},~{}y^{(k)})\}^{\infty}_{k=0}$ generated by (13) converges to $(x^{*},~{}Q^{-1}|x^{*}|)$ and $x^{*}$ is the unique solution of GAVEs (1).
@@
The corollary will be proved if we can show that (46) and (47) imply (26) and (27). The proof is divided into two cases.
@@
    We first consider $\dfrac{2c\beta}{(1-a)(1-d)-c\alpha}<\theta\leq 1$. It follows from (46) that
@@
    $\dfrac{c\beta-(a-1)}{a(d-1)}<0<\theta,$ (48)
    $\theta\leq 1<\dfrac{c\beta-(a+1)}{a(d-1)}.$ (49)
    From the inequality (48), we have
@@
    $a(1-\theta)+\theta ad-c\beta<1.$ (50)
    From the inequality (49), we have
@@
    $a(1-\theta)+\theta ad-c\beta>-1.$ (51)
    Note that $0<\theta\leq 1$, combining (50) and (51), we have
@@
    which is the inequality (26).
    On the other hand, it follows from (46) that $c\alpha<(1-a)(1-d)$ (otherwise, we get $02c\beta<$ 0, a contradiction). Multiplying both sides of $\dfrac{2c\beta}{(1-a)(1-d)-c\alpha}<\theta$ by $(1-a)(1-d)-c\alpha$, we have
    $2c\beta<\theta(1-a)(1-d)-\theta c\alpha.$ (52)
    From (52), we infer that (remind that $0<\theta\leq 1$)
@@
    Now we turn to $1<\theta<\dfrac{2(1-a)-2c\beta}{(1-a)(1+d)+c\alpha}$. On the condition of (46) and (47), we have
    $\dfrac{c\beta+a-1}{a(d+1)}<1$ $<\theta,$ (53)
    $\theta<\dfrac{2(1-a)-2c\beta}{(1-a)(1+d)+c\alpha}$ $<\dfrac{c\beta+a+1}{a(d+1)}.$ (54)
    From (53), we have
@@
    $a(\theta-1)+\theta ad-c\beta>-1.$ (55)
    From (54), we have
@@
    $a(\theta-1)+\theta ad-c\beta<1.$ (56)
    Note that $\theta>1$, combining (55) and (56), we have
@@
    which is the inequality (26).
    On the other hand, if $a<1$, we have $(1-a)(1+d)+c\alpha>0$. According to the first inequality of (54), we have
@@
In the following, we will apply the results of the Theorem 1 to some existing methods.
@@
$\tau_{\rm RMS}\|M_{\rm RMS}^{-1}B\|<(1-\|M_{\rm RMS}^{-1}N_{\rm RMS}\|)(1-|1-\tau_{\rm RMS}|),$ (58)
then GAVEs (1) has a unique solution $x^{*}$ for any $b\in\mathbb{R}^{n}$ and the sequence $\{x^{(k)}\}_{k=0}^{\infty}$ generated by the RMS iteration (3) converges to $x^{*}$.
Remark 2.
The conditions (57) and (58) appear in [27, Lemma 3]. However, the existence and uniqueness of the solution to GAVEs (1) is not considered there.
@@
$\|{(M_{\rm NMS}+\Omega_{\rm NMS})}^{-1}(N_{\rm NMS}+\Omega_{\rm NMS})\|+\|{(M_{\rm NMS}+\Omega_{\rm NMS})}^{-1}B\|<1,$ (59)
then GAVEs (1) has a unique solution $x^{*}$ for any $b\in\mathbb{R}^{n}$ and the sequence $\{x^{(k)}\}_{k=0}^{\infty}$ generated by the NMS iteration [34] converges to $x^{*}$.
Remark 3.
In [34, Corollary 4.2.], the authors show that the sequence $\{x^{(k)}\}_{k=0}^{\infty}$ generated by the NMS iteration converges to a solution of GAVEs (1) if
$\|(M_{\rm NMS}+\Omega_{\rm NMS})^{-1}\|(\|N_{\rm NMS}+\Omega_{\rm NMS}\|+\|B\|)<1,$ (60)
which implies (59) while the reverse is generally not true. Moreover, the existence and uniqueness of the solution to GAVEs (1) is not considered in [34, Corollary 4.2.].
Corollary 4.
@@
$\|(A+\Omega_{\rm MN})^{-1}\Omega_{\rm MN}\|+\|(A+\Omega_{\rm MN})^{-1}B\|<1,$ (61)
then GAVEs (1) has a unique solution $x^{*}$ for any $b\in\mathbb{R}^{n}$ and the sequence $\{x^{(k)}\}_{k=0}^{\infty}$ generated by the MN iteration [28] converges to $x^{*}$.
Remark 4.
In [28, Theorem 3.1], the authors show that the sequence $\{x^{(k)}\}_{k=0}^{\infty}$ generated by the MN iteration converges to a solution of GAVEs (1) if
$\|(A+\Omega_{\rm MN})^{-1}\|(\|\Omega_{\rm MN}\|+\|B\|)<1.$ (62)
It is obvious that (62) implies (61) while the reverse is generally not true. In addition, the existence and uniqueness of the solution to GAVEs (1) is not considered in [28, Theorem 3.1].
Corollary 5.
Assume that $A,B\in\mathbb{R}^{n\times n}$ are given and $A$ is nonsingular. If $\|A^{-1}B\|<1$, then GAVEs (1) has a unique solution $x^{*}$ for any $b\in\mathbb{R}^{n}$ and the sequence $\{x^{(k)}\}_{k=0}^{\infty}$ generated by the Picard iteration [26] converges to $x^{*}$.
Remark 5.
@@
$|1-\alpha_{\rm MGSOR}||1-\beta_{\rm MGSOR}|<1$ (63a)
@@
$\alpha_{\rm MGSOR}\beta_{\rm MGSOR}\|A^{-1}BQ_{\rm MGSOR}\|\|Q_{\rm MGSOR}^{-1}\|<(1-|1-\alpha_{\rm MGSOR}|)(1-|1-\beta_{\rm MGSOR}|),$ (63b)
then GAVEs (1) has a unique solution $x^{*}$ for any $b\in\mathbb{R}^{n}$ and the sequence $\{(x^{(k)},y^{(k)})\}_{k=0}^{\infty}$ generated by the MGSOR iteration (19) converges to $(x^{*},y^{*}=Q_{\rm MGSOR}^{-1}|x^{*}|)$ for any given initial vectors $x^{(0)},y^{(0)}\in\mathbb{R}^{n}$.
When $B=I$ we can have the following Corollary 7 directly from Corollary 6.
Corollary 7.
@@
$|1-\alpha_{\rm MGSOR}||1-\beta_{\rm MGSOR}|<1$ (64a)
@@
$\alpha_{\rm MGSOR}\beta_{\rm MGSOR}\|A^{-1}Q_{\rm MGSOR}\|\|Q_{\rm MGSOR}^{-1}\|<(1-|1-\alpha_{\rm MGSOR}|)(1-|1-\beta_{\rm MGSOR}|),$ (64b)
then AVEs (2) has a unique solution $x^{*}$ for any $b\in\mathbb{R}^{n}$ and the sequence $\{(x^{(k)},y^{(k)})\}_{k=0}^{\infty}$ generated by the MGSOR iteration (15) converges to $(x^{*},y^{*}=Q_{\rm MGSOR}^{-1}|x^{*}|)$ for any given initial vectors $x^{(0)},y^{(0)}\in\mathbb{R}^{n}$.
Remark 6.
The convergence condition (64) is new, which is not appeared in [32]. In addition, the existence and uniqueness of the solution to AVEs (2) is not considered in [32].
When $\alpha_{\rm MGSOR}=1$, $\beta_{\rm MGSOR}=\tau_{\rm MFPI}$ and $Q_{\rm MGSOR}=Q_{\rm MFPI}$, the following Corollary 8 follows from Corollary 6.
Corollary 8.
@@
then GAVEs (1) has a unique solution $x^{*}$ for any $b\in\mathbb{R}^{n}$ and the sequence $\{(x^{(k)},y^{(k)})\}_{k=0}^{\infty}$ generated by the MFPI iteration (21) converges to $(x^{*},y^{*}=Q_{\rm MFPI}^{-1}|x^{*}|)$ for any given initial vectors $x^{(0)},y^{(0)}\in\mathbb{R}^{n}$.
When $\alpha_{\rm MGSOR}=1$, $\beta_{\rm MGSOR}=\tau_{\rm MFPI}$, $Q_{\rm MGSOR}=Q_{\rm MFPI}$ and $B=I$ we can establish the following Corollary 9 from Corollary 7.
Corollary 9.
@@
$0<\|A^{-1}Q_{\rm MFPI}\|\|Q^{-1}_{\rm MFPI}\|<1\quad\text{and}\quad 0<\tau_{\rm MFPI}<\frac{2}{1+\|A^{-1}Q_{\rm MFPI}\|\|Q^{-1}_{\rm MFPI}\|},$ (65)
then AVEs (2) has a unique solution $x^{*}$ for any $b\in\mathbb{R}^{n}$ and the sequence $\{(x^{(k)},y^{(k)})\}_{k=0}^{\infty}$ generated by the MFPI iteration (10) converges to $(x^{*},y^{*}=Q_{\rm MFPI}^{-1}|x^{*}|)$ for any given initial vectors $x^{(0)},y^{(0)}\in\mathbb{R}^{n}$.
Remark 7.
In [30], the authors show that the sequence $\{x^{(k)}\}_{k=0}^{\infty}$ generated by the MFPI iteration (10) converges to the unique solution $x^{*}$ of AVEs (2) if
$0<\nu<\dfrac{1}{\ell\sqrt{1+\varrho^{2}}}\quad\text{and}\quad\dfrac{1-\sqrt{1-\ell^{2}\nu^{2}}}{1-\ell\varrho\nu}<\tau_{\rm MFPI}<\dfrac{1+\sqrt{1-\ell^{2}\nu^{2}}}{1+\ell\varrho\nu},$ (66)
@@
In addition, $0<\nu<\dfrac{1}{\ell\sqrt{1+\varrho^{2}}}$ implies $\dfrac{1-\sqrt{1-\ell^{2}\nu^{2}}}{1-\ell\varrho\nu}>0$ and $\frac{2}{1+\|A^{-1}Q_{\rm MFPI}\|\|Q^{-1}_{\rm MFPI}\|}\geq\frac{2}{1+\|A^{-1}\|\|Q_{\rm MFPI}\|\|Q^{-1}_{\rm MFPI}\|}>\dfrac{1+\sqrt{1-\ell^{2}\nu^{2}}}{1+\ell\varrho\nu}$. Hence, (65) is weaker than (66).
If $Q_{\rm MFPI}=I$ and $\tau_{\rm MFPI}=\tau_{\rm FPI}$, the following corollary follows from Corollary 8.
Corollary 10.
@@
then GAVEs (1) has a unique solution $x^{*}$ for any $b\in\mathbb{R}^{n}$ and the sequence $\{(x^{(k)},y^{(k)})\}_{k=0}^{\infty}$ generated by the FPI iteration (18) converges to $(x^{*},y^{*}=|x^{*}|)$ for any given initial vectors $x^{(0)},y^{(0)}\in\mathbb{R}^{n}$.
When $Q_{\rm MFPI}=I$, $\tau_{\rm MFPI}=\tau_{\rm FPI}$ and $B=I$ we can have the following result follows from Corollary 9.
Corollary 11.
@@
$0<\nu<1\quad\text{and}\quad 0<\tau_{\rm FPI}<\frac{2}{1+\nu},$ (67)
then AVEs (2) has a unique solution $x^{*}$ for any $b\in\mathbb{R}^{n}$ and the sequence $\{(x^{(k)},y^{(k)})\}_{k=0}^{\infty}$ generated by the FPI iteration (8) converges to $(x^{*},y^{*}=|x^{*}|)$ for any given initial vectors $x^{(0)},y^{(0)}\in\mathbb{R}^{n}$.
Remark 8.
In [11], Ke show that the sequence $\{x^{(k)}\}_{k=0}^{\infty}$ generated by the FPI iteration (8) converges to the unique solution of AVEs (2) if
$0<\nu<\frac{\sqrt{2}}{2}\quad\text{and}\quad\frac{1-\sqrt{1-\nu^{2}}}{1-\nu}<\tau_{\rm FPI}<\frac{1+\sqrt{1-\nu^{2}}}{1+\nu}.$ (68)
It is obvious that $0<\dfrac{1-\sqrt{1-\nu^{2}}}{1-\nu}$ and $\dfrac{1+\sqrt{1-\nu^{2}}}{1+\nu}<\dfrac{2}{1+\nu}$ when $10<\nu<$ 1. Hence, $\{\nu:0<\nu<\frac{\sqrt{2}}{2}\}\subseteq\{\nu:0<\nu<1\}$ and $\{\tau_{\rm FPI}:\frac{1-\sqrt{1-\nu^{2}}}{1-\nu}<\tau_{\rm FPI}<\frac{1+\sqrt{1-\nu^{2}}}{1+\nu}\}\subseteq\{\tau_{\rm FPI}:0<\tau_{\rm FPI}<\frac{2}{1+\nu}\}$. In other words, (67) is weaker than (68). Moreover, we get the theoretical improvement here without using the modification strategy proposed in [30].
When $\alpha_{\rm MGSOR}=\beta_{\rm MGSOR}=\omega_{\rm SOR}$ and $Q_{\rm MGSOR}=I$, the following corollary follows from Corollary 6.
Corollary 12.
@@
$0<\omega_{\rm SOR}<2\quad\text{and}\quad 0<\|A^{-1}B\|<\Big{(}\dfrac{1-|1-\omega_{\rm SOR}|}{\omega_{\rm SOR}}\Big{)}^{2},$
then GAVEs (1) has a unique solution $x^{*}$ for any $b\in\mathbb{R}^{n}$ and the sequence $\{(x^{(k)},y^{(k)})\}_{k=0}^{\infty}$ generated by the SOR-like iteration (20) converges to $(x^{*},y^{*}=|x^{*}|)$ for any given initial vectors $x^{(0)},y^{(0)}\in\mathbb{R}^{n}$.
When $\alpha_{\rm MGSOR}=\beta_{\rm MGSOR}=\omega_{\rm SOR}$, $Q_{\rm MGSOR}=I$ and $B=I$ we can have the following result from Corollary 7.
Corollary 13.
@@
$0<\nu<1\quad\text{and}\quad 0<\omega_{\rm SOR}<\frac{2-2\sqrt{\nu}}{1-\nu},$ (69)
then AVEs (2) has a unique solution $x^{*}$ for any $b\in\mathbb{R}^{n}$ and the sequence $\{(x^{(k)},y^{(k)})\}_{k=0}^{\infty}$ generated by the SOR-like iteration (7) converges to $(x^{*},y^{*}=|x^{*}|)$ for any given initial vectors $x^{(0)},y^{(0)}\in\mathbb{R}^{n}$.
Remark 9.
In [10, Corollary 3.1], under a kernel norm, Ke and Ma show that the SOR-like iteration converges to the unique solution of GAVEs (1) if
$0<\nu<1\quad\text{and}\quad 1-\frac{2}{3+\sqrt{5}}<\omega_{\rm SOR}<\min\left\{1+\frac{2}{3+\sqrt{5}},\sqrt{\frac{2}{(3+\sqrt{5})\nu}}\right\}.$ (70)
If $0<\nu<\left(\frac{1+\sqrt{5}}{5+\sqrt{5}}\right)^{2}$, we have $\frac{2-2\sqrt{\nu}}{1-\nu}>1+\frac{2}{3+\sqrt{5}}\geq\min\left\{1+\frac{2}{3+\sqrt{5}},\sqrt{\frac{2}{(3+\sqrt{5})\nu}}\right\}$, which implies that (69) is weaker than (70).